# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ Abstract
Diffusion models have recently shown considerable potential in solving Bayesian inverse problems when used as priors. However, sampling from the resulting denoising posterior distributions remains a challenge as it involves intractable terms. To tackle this issue, state-of-the-art approaches formulate the problem as that of sampling from a surrogate diffusion model targeting the posterior and decompose its scores into two terms: the prior score and an intractable guidance term. While the former is replaced by the pre-trained score of the considered diffusion model, the guidance term has to be estimated. In this paper, we propose a novel approach that utilises a decomposition of the transitions which, in contrast to previous methods, allows a trade-off between the complexity of the intractable guidance term and that of the prior transitions. We validate the proposed approach through extensive experiments on linear and nonlinear inverse problems, including challenging cases with latent diffusion models as priors, and demonstrate its effectiveness in reconstructing electrocardiogram (ECG) from partial measurements for accurate cardiac diagnosis.
@@ 1 Introduction
Inverse problems aim to reconstruct signals from incomplete and noisy observations and are prevalent across various fields. In signal and image processing, common examples include signal deconvolution, image restoration, and tomographic image reconstruction Stuart (2010); Idier (2013). Other applications extend to protein backbone motif scaffolding Watson et al. (2023) and urban mobility modeling Jiang et al. (2023). Due to the presence of noise and the inherent complexity of the measurement process, these problems are typically ill-posed, meaning they have an infinite number of possible solutions. While many of these solutions may fit the observed data, only a few align with the true underlying signal. Bayesian inverse problems provide a principled framework for addressing the challenges of signal reconstruction by incorporating prior knowledge, enabling more plausible and meaningful solutions. The a priori knowledge about the signal $x$ to be reconstructed is captured in the prior distribution $q(x)$, while the information about the observation $y$ is encoded in the likelihood function $p(y|x)$. Given these components, solving the inverse problem boils down to sampling from the posterior distribution $\pi(x)\propto p(y|x)q(x)$, which integrates both prior knowledge and observational data.
The choice of the prior distribution is crucial for achieving accurate reconstructions. If the goal is to reconstruct high-resolution data, the prior must be able to model data with similar fine detail and complexity. Recently, denoising diffusion models (DDM) Sohl-Dickstein et al. (2015); Song & Ermon (2019); Ho et al. (2020); Song et al. (2021b) have emerged as a powerful approach in this context. These models can generate highly realistic and detailed reconstructions and are becoming increasingly popular as priors. Using a DDM, a sample $X_{0}$ being approximately distributed according to a given data distribution $q$ of interest can be generated by iteratively denoising an initial sample $X_{n}$ from a standard Gaussian distribution. The denoising process $(X_{k})_{k=n}^{0}$ consists of the iterative refinement of $X_{k}$ for $k\in\llbracket 1,n\rrbracket$ based on a parametric approximation of the score $\nabla\log q_{k}$, where $q_{k}(x_{k})=\int q(x_{0})q_{k|0}(x_{k}|x_{0}),\mathrm{d}x_{0}$, with $q_{k|0}(\cdot|x_{0})$ being a Gaussian noising transition.
@@
The problem of sampling from the posterior in Bayesian inverse problems with DDM priors has recently been addressed in several papers Kadkhodaie & Simoncelli (2020); Song et al. (2021b); Kawar et al. (2021; 2022); Lugmayr et al. (2022); Ho et al. (2022); Chung et al. (2023); Song et al. (2023a). These methods typically modify the denoising process to account for the observation $y$. One principled approach for accurate posterior sampling involves skewing the denoising process at each stage $k$ using the score $\nabla\log\pi_{k}$, where $\pi_{k}$ is defined analogously to $q_{k}$, with $q$ replaced by the posterior $\pi$. This guides the denoising process in a manner that is consistent with the observation. Notably, the posterior score $\nabla\log\pi_{k}$ decomposes into two components: the prior score and an additional intractable term commonly referred to as guidance. Previous works leverage pre-trained DDMs for the prior score, while providing various training-free approximations for the guidance term Ho et al. (2022); Chung et al. (2023); Song et al. (2023a); Finzi et al. (2023). This framework enables solving a wide range of inverse problems for a given prior. However, despite the notable success of these methods, the approximations often introduce significant errors, leading to inaccurate reconstructions when the posterior distribution is highly multimodal, the measurement process is strongly nonlinear, or the data is heavily contaminated by noise.
@@
We begin with the observation that the posterior denoising step, which transforms a sample $X_{k+1}$ into a sample $X_{k}$, does not necessarily require the conditional scores $\nabla\log\pi_{k+1}$ or $\nabla\log\pi_{k}$. Instead, we demonstrate that this step can be decomposed into two intermediate phases: first, denoising $X_{k+1}$ into an intermediate midpoint state $X_{\ell_{k}}$, where ${\ell_{k}}<k$, and then noising back to obtain $X_{k}$, unconditionally on the observation $y$. This decomposition introduces an additional degree of freedom, as it only requires the estimation of the guidance term at the intermediate step ${\ell_{k}}$ rather than step $k+1$. Building on this insight, we introduce Midpoint Guidance Posterior Sampling  (MGPS), a novel diffusion posterior sampling scheme that explicitly leverages this approach. Our algorithm develops a principled approximation of the denoising transition by utilizing a Gaussian variational approximation, combined with the guidance approximation proposed by Chung et al. (2023) at the intermediate steps ${\ell_{k}}$. The strong empirical performance of our method is demonstrated through an extensive set of experiments. In particular, we validate our approach on a Gaussian mixture toy example, as well as various linear and nonlinear image reconstruction tasks–inpainting, super-resolution, phase retrieval, deblurring, JPEG dequantization, high-dynamic range–using both pixel-space and latent diffusion models (LDM). We also apply our posterior sampling method to imputation tasks for 12-lead electrocardiograms (ECGs). Our results demonstrate that training-free posterior sampling performs on par with diffusion models specifically trained for ECG imputation Alcaraz & Strodthoff (2022). Moreover, our method demonstrates significantly improved performance compared to existing approaches.
@@ 2 Posterior sampling with DDM prior
$\pi(x)\vcentcolon=g(x)q(x)/\mathcal{Z}\,,$ (2.1)
where $g:\mathbb{R}^{{d}}\to\mathbb{R}$ is a non-negative and differentiable potential that can be evaluated pointwise, $q$ is a prior distribution, and $\smash{\mathcal{Z}\vcentcolon=\int g(x)q(x)\,\mathrm{d}x}$ is the normalizing constant. We are interested in the case where a DDM $p^{\theta}_{0}$ for the prior has been pre-trained for the prior based on i.i.d. observations from $q$. As a by-product, we also have parametric approximations $(s^{\theta}_{k})_{k=1}^{n}$ of the scores $(\nabla\log q_{k})_{k=1}^{n}$, where the marginals $(q_{k})_{k=1}^{n}$ are defined as $q_{k}(x_{k})\vcentcolon=\int q(x_{0})\,q_{k|0}(x_{k}|x_{0})\,\mathrm{d}x_{0}$ with $q_{k|0}(x_{k}|x_{0})\vcentcolon=\mathcal{N}(x_{k};\sqrt{\alpha_{k}}x_{0},v_{k}\mathrm{I}_{d})$ and $v_{k}\vcentcolon=1-\alpha_{k}$. Typically, the sequence $(\alpha_{k})_{k=0}^{n}$ is a decreasing sequence with $\alpha_{0}=1$ and $\alpha_{n}$ approximately equals zero for $n$ large enough.
These score approximations enable the generation of new samples from $q$ according to one of the DDM sampling schemes Song & Ermon (2019); Ho et al. (2020); Song et al. (2021a; b); Karras et al. (2022). All these approaches boil down to simulating a Markov chain $(X_{k})_{k=n}^{0}$ backwards in time starting from a standard Gaussian $p_{n}\vcentcolon=\mathcal{N}(\mathrm{0},\mathrm{I}_{{d}})$ and following the Gaussian transitions $p^{\theta}_{k|k+1}(x_{k}|x_{k+1})=\mathcal{N}(x_{k};\smash{m^{\theta}_{k|k+1}}(x_{k+1}),v_{k|k+1}\mathrm{I}_{{d}})$, where the mean functions $\smash{m^{\theta}_{k|k+1}}$ can be derived from the learned DDM, and $v_{k|k+1}>0$ are fixed and pre-defined variances. The backward transitions are designed in such a way that the marginal law of $X_{k}$ approximates $q_{k}$; see Section A.1 for more details on the form of the backward transitions.
In our problem setup, we assume we only have access to the approximated scores of the prior and no observation from neither posterior $\pi$ nor the prior $q$. This setup encompasses of solving an inverse problem without training a conditional model from scratch based on a paired signal/observation dataset; a requirement typically imposed by conditional DDM frameworks Song et al. (2021b); Batzolis et al. (2021); Tashiro et al. (2021); Saharia et al. (2022). On the other hand, we require access to $g$. This setup includes many applications in Bayesian inverse problems, e.g., image restoration, motif scaffolding in protein design Trippe et al. (2023); Wu et al. (2023), and trajectory control in traffic-simulation frameworks Jiang et al. (2023).
@@
Since we have access to a DDM model for $q$, a natural approach to sampling from $\pi$ is to define a DDM approximation based on the pre-trained score approximations for $q_{k}$. Following the DDM approach, the basic idea here is to sample sequentially from the marginals $\pi_{k}(x_{k})\vcentcolon=\int\pi(x_{0})q_{k|0}(x_{k}|x_{0})\,\mathrm{d}x_{0}$ by relying on approximations of the conditional scores $(\nabla\log\pi_{k})_{k=1}^{n}$. The latter can be expressed in terms of the unconditional scores by using
$\pi_{k}(x_{k})\propto\int g(x_{0})q(x_{0})q_{k|0}(x_{k}|x_{0})\,\mathrm{d}x_{0}\,,$ (2.2)
where after defining the backward transition kernel $q_{0|k}(x_{0}|x_{k})\vcentcolon=q(x_{0})q_{k|0}(x_{k}|x_{0})\big{/}q_{k}(x_{k})$ yields
$\pi_{k}(x_{k})\propto g_{k}(x_{k})q_{k}(x_{k})\,,\quad\text{where}\quad g_{k}(x_{k})\vcentcolon=\int g(x_{0})q_{0|k}(x_{0}|x_{k})\,\mathrm{d}x_{0}\,.$ (2.3)
@@
$\nabla\log\pi_{k}(x_{k})=\nabla\log q_{k}(x_{k})+\nabla\log g_{k}(x_{k})\,.$
While we have a parametric approximation of the first term on the r.h.s., the bottleneck is the second term, which is intractable due to the integration of $g$ against the conditional distribution $\smash{q_{0|k}(\cdot|x_{k})}$.
@@
To circumvent this computational bottleneck, previous works have proposed approximations that involve replacing $q_{0|k}(\cdot|x_{k})$ in (2.3) with either a Dirac delta mass Chung et al. (2023) or a Gaussian approximation Ho et al. (2022); Song et al. (2023a)
$\smash{q_{0|k}(x_{0}|x_{k})\approx\mathcal{N}(x_{0};\smash{m^{\theta}_{0|k}}(x_{k}),v_{0|k}\mathrm{I}_{d})}\,,$ (2.4)
@@
where $\smash{m^{\theta}_{0|k}}$ is a parametric approximation of $\smash{\smash{m_{0|k}}(x_{k})\vcentcolon=\int x_{0}\,q_{0|k}(x_{0}|x_{k})\,\mathrm{d}x_{0}}$ and $v_{0|k}$ is a tuning parameter. Using Tweedie’s formula Robbins (1956), it holds that $\smash{m_{0|k}}(x_{k})=(-x_{k}+\sqrt{\alpha_{k}}\,\nabla\log q_{k}(x_{k}))/v_{k}$, which suggests the approximation $\smash{m^{\theta}_{0|k}}(x_{k})\vcentcolon=(-x_{k}+\sqrt{\alpha_{k}}s^{\theta}_{k}(x_{k}))/v_{k}\,.$ As for the variance parameter $v_{0|k}$, Ho et al. (2022) suggest setting $v_{0|k}=v_{k}/\alpha_{k}$, while Song et al. (2023a) use $v_{0|k}=v_{k}$. When $g$ is the likelihood of a linear model, i.e., $g(x)\vcentcolon=\mathcal{N}(y;\mathrm{A}x,\sigma_{y}^{2}\mathrm{I}_{d_{y}})$, for $\mathrm{A}\in\mathbb{R}^{{d_{y}}\times{d}}$ and $\sigma_{y}>0$, it can be exactly integrated against the Gaussian approximation (2.4), yielding $g_{k}(x_{k})\approx\mathcal{N}(y;\mathrm{A}\smash{m^{\theta}_{0|k}}(x_{k}),\sigma_{y}^{2}\mathrm{I}_{d_{y}}+v_{0|k}\mathrm{A}\mathrm{A}^{\intercal})$. Chung et al. (2023), on the other hand, use the pointwise approximation, yielding $g_{k}(x_{k})\approx g(\smash{m^{\theta}_{0|k}}(x_{k}))$. We denote by $\tilde{g}_{k}$ the resulting approximation stemming from any of these methods. Approximate samples from the posterior distribution are then drawn by simulating a backward Markov chain $(X_{k})_{k=n}^{0}$, where $X_{n}\sim p_{n}$ and then, given $X_{k+1}$, $X_{k}$ is obtained via the update
@@
$X_{k}\vcentcolon=\tilde{X}_{k}+w_{k+1}(X_{k+1})\nabla\log\tilde{g}_{k+1}(X_{k+1})\,,\quad\text{where}\quad\tilde{X}_{k}\sim p^{\theta}_{k|k+1}(\cdot|X_{k+1})\,,$ (2.5)
@@
and $w_{k+1}$ is a method-dependent weighting function. For instance, Chung et al. (2023) use $w_{k+1}(X_{k+1})=\zeta/\|y-\mathrm{A}\smash{m^{\theta}_{0|k+1}}(X_{k+1})\|$, with $\zeta\in[0,1]$. We note that the update (2.5) in general involves a vector-Jacobian product, e.g., considering the approximation suggested by Chung et al. (2023), $\nabla_{x_{k}}\log\tilde{g}_{k}(x_{k})=\nabla_{x_{k}}\smash{m^{\theta}_{0|k}}(x_{k})^{\intercal}\nabla_{x_{0}}\left.\log g(x_{0})\right|_{x_{0}=\smash{m^{\theta}_{0|k}}(x_{k})}$, which incurs an additional computational cost compared to an unconditional diffusion step.

### v2

@@ Abstract
Diffusion models have recently shown considerable potential in solving Bayesian inverse problems when used as priors. However, sampling from the resulting denoising posterior distributions remains a challenge as it involves intractable terms. To tackle this issue, state-of-the-art approaches formulate the problem as that of sampling from a surrogate diffusion model targeting the posterior and decompose its scores into two terms: the prior score and an intractable guidance term. While the former is replaced by the pre-trained score of the considered diffusion model, the guidance term has to be estimated. In this paper, we propose a novel approach that utilises a decomposition of the transitions which, in contrast to previous methods, allows a trade-off between the complexity of the intractable guidance term and that of the prior transitions. We validate the proposed approach through extensive experiments on linear and nonlinear inverse problems, including challenging cases with latent diffusion models as priors. We then demonstrate its applicability to various modalities and its promising impact on public health by tackling cardiovascular disease diagnosis through the reconstruction of incomplete electrocardiograms. The code is publicly available at https://github.com/yazidjanati/mgps.
@@ 1 Introduction
Inverse problems aim to reconstruct signals from incomplete and noisy observations and are prevalent across various fields. In signal and image processing, common examples include signal deconvolution, image restoration, and tomographic image reconstruction Stuart (2010); Idier (2013). Other applications extend to protein backbone motif scaffolding Watson et al. (2023) and urban mobility modeling Jiang et al. (2023). Due to the presence of noise and the inherent complexity of the measurement process, these problems are typically ill-posed, meaning they have an infinite number of possible solutions. While many of these solutions may fit the observed data, only a few align with the true underlying signal. Bayesian inverse problems provide a principled framework for addressing the challenges of signal reconstruction by incorporating prior knowledge, enabling more plausible and meaningful solutions. The a priori knowledge about the signal $\boldsymbol{x}$ to be reconstructed is captured in the prior distribution $q(\boldsymbol{x})$, while the information about the observation ${\boldsymbol{y}}$ is encoded in the likelihood function $p({\boldsymbol{y}}|\boldsymbol{x})$. Given these components, solving the inverse problem boils down to sampling from the posterior distribution $\pi(\boldsymbol{x})\propto p({\boldsymbol{y}}|\boldsymbol{x})q(\boldsymbol{x})$, which integrates both prior knowledge and observational data.
The choice of the prior distribution is crucial for achieving accurate reconstructions. If the goal is to reconstruct high-resolution data, the prior must be able to model data with similar fine detail and complexity. Recently, denoising diffusion models (DDM) Sohl-Dickstein et al. (2015); Song & Ermon (2019); Ho et al. (2020); Song et al. (2021b) have emerged as a powerful approach in this context. These models can generate highly realistic and detailed reconstructions and are becoming increasingly popular as priors. Using a DDM, a sample $\mathbf{X}_{0}$ being approximately distributed according to a given data distribution $q$ of interest can be generated by iteratively denoising an initial sample $\mathbf{X}_{n}$ from a standard Gaussian distribution. The denoising process $(\mathbf{X}_{k})_{k=n}^{0}$ consists of the iterative refinement of $\mathbf{X}_{k}$ for $k\in\llbracket 1,n\rrbracket$ based on a parametric approximation of the score $\nabla\log q_{k}$, where $q_{k}(\boldsymbol{x}_{k})=\int q(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0}),\mathrm{d}\boldsymbol{x}_{0}$, with $q_{k|0}(\cdot|\boldsymbol{x}_{0})$ being a Gaussian noising transition.
@@
The problem of sampling from the posterior in Bayesian inverse problems with DDM priors has recently been addressed in several papers Kadkhodaie & Simoncelli (2020); Song et al. (2021b); Kawar et al. (2021; 2022); Lugmayr et al. (2022); Ho et al. (2022); Chung et al. (2023); Song et al. (2023a). These methods typically modify the denoising process to account for the observation ${\boldsymbol{y}}$. One principled approach for accurate posterior sampling involves skewing the denoising process at each stage $k$ using the score $\nabla\log\pi_{k}$, where $\pi_{k}$ is defined analogously to $q_{k}$, with $q$ replaced by the posterior $\pi$. This guides the denoising process in a manner that is consistent with the observation. Notably, the posterior score $\nabla\log\pi_{k}$ decomposes into two components: the prior score and an additional intractable term commonly referred to as guidance. Previous works leverage pre-trained DDMs for the prior score, while providing various training-free approximations for the guidance term Ho et al. (2022); Chung et al. (2023); Song et al. (2023a); Finzi et al. (2023). This framework enables solving a wide range of inverse problems for a given prior. However, despite the notable success of these methods, the approximations often introduce significant errors, leading to inaccurate reconstructions when the posterior distribution is highly multimodal, the measurement process is strongly nonlinear, or the data is heavily contaminated by noise.
@@
We begin with the observation that the posterior denoising step, which transforms a sample $\mathbf{X}_{k+1}$ into a sample $\mathbf{X}_{k}$, does not necessarily require the conditional scores $\nabla\log\pi_{k+1}$ or $\nabla\log\pi_{k}$. Instead, we demonstrate that this step can be decomposed into two intermediate phases: first, denoising $\mathbf{X}_{k+1}$ into an intermediate midpoint state $\mathbf{X}_{\ell_{k}}$, where ${\ell_{k}}<k$, and then noising back to obtain $\mathbf{X}_{k}$, unconditionally on the observation ${\boldsymbol{y}}$. This decomposition introduces an additional degree of freedom, as it only requires the estimation of the guidance term at the intermediate step ${\ell_{k}}$ rather than step $k+1$. Building on this insight, we introduce Midpoint Guidance Posterior Sampling  (MGPS), a novel diffusion posterior sampling scheme that explicitly leverages this approach. Our algorithm develops a principled approximation of the denoising transition by utilizing a Gaussian variational approximation, combined with the guidance approximation proposed by Chung et al. (2023) at the intermediate steps ${\ell_{k}}$. The strong empirical performance of our method is demonstrated through an extensive set of experiments. In particular, we validate our approach on a Gaussian mixture toy example, as well as various linear and nonlinear image reconstruction tasks–inpainting, super-resolution, phase retrieval, deblurring, JPEG dequantization, high-dynamic range–using both pixel-space and latent diffusion models (LDM). Finally, we demonstrate the versatility of our approach by applying it to cardiovascular diagnosis using reconstructed electrocardiograms, showing that MGPS achieves significant improvements over competing methods.
@@ 2 Posterior sampling with DDM prior
$\pi(\boldsymbol{x})\vcentcolon=p({\boldsymbol{y}}|\boldsymbol{x})q(\boldsymbol{x})/\mathcal{Z}\,,$ (2.1)
where $p({\boldsymbol{y}}|\cdot):\mathbb{R}^{{d}}\to\mathbb{R}$ is a non-negative and differentiable likelihood that can be evaluated pointwise, $q$ is a prior distribution, and $\smash{\mathcal{Z}\vcentcolon=\int p({\boldsymbol{y}}|\boldsymbol{x})q(\boldsymbol{x})\,\mathrm{d}\boldsymbol{x}}$ is the normalizing constant. We are interested in the case where a DDM $p^{\theta}_{0}$ for the prior has been pre-trained for the prior based on i.i.d. observations from $q$. As a by-product, we also have parametric approximations $(\boldsymbol{s}^{\theta}_{k})_{k=1}^{n}$ of the scores $(\nabla\log q_{k})_{k=1}^{n}$, where the marginals $(q_{k})_{k=1}^{n}$ are defined as $q_{k}(\boldsymbol{x}_{k})\vcentcolon=\int q(\boldsymbol{x}_{0})\,q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\,\mathrm{d}\boldsymbol{x}_{0}$ with $q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\vcentcolon=\mathcal{N}(\boldsymbol{x}_{k};\sqrt{\alpha_{k}}\boldsymbol{x}_{0},v_{k}\mathbf{I}_{d})$ and $v_{k}\vcentcolon=1-\alpha_{k}$. Typically, the sequence $(\alpha_{k})_{k=0}^{n}$ is a decreasing sequence with $\alpha_{0}=1$ and $\alpha_{n}$ approximately equals zero for $n$ large enough.
These score approximations enable the generation of new samples from $q$ according to one of the DDM sampling schemes Song & Ermon (2019); Ho et al. (2020); Song et al. (2021a; b); Karras et al. (2022). All these approaches boil down to simulating a Markov chain $(\mathbf{X}_{k})_{k=n}^{0}$ backwards in time starting from a standard Gaussian $p_{n}\vcentcolon=\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{{d}})$ and following the Gaussian transitions $p^{\theta}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})=\mathcal{N}(\boldsymbol{x}_{k};\smash{\boldsymbol{m}^{\theta}_{k|k+1}}(\boldsymbol{x}_{k+1}),v_{k|k+1}\mathbf{I}_{{d}})$, where the mean functions $\smash{\boldsymbol{m}^{\theta}_{k|k+1}}$ can be derived from the learned DDM, and $v_{k|k+1}>0$ are fixed and pre-defined variances. The backward transitions are designed in such a way that the marginal law of $\mathbf{X}_{k}$ approximates $q_{k}$; see Section A.1 for more details on the form of the backward transitions.
In our problem setup, we assume we only have access to the approximated scores of the prior and no observation from neither posterior $\pi$ nor the prior $q$. This setup encompasses of solving an inverse problem without training a conditional model from scratch based on a paired signal/observation dataset; a requirement typically imposed by conditional DDM frameworks Song et al. (2021b); Batzolis et al. (2021); Tashiro et al. (2021); Saharia et al. (2022). On the other hand, we require access to $p({\boldsymbol{y}}|\cdot)$. This setup includes many applications in Bayesian inverse problems, e.g., image restoration, motif scaffolding in protein design Trippe et al. (2023); Wu et al. (2023), and trajectory control in traffic-simulation frameworks Jiang et al. (2023).
@@
Since we have access to a DDM model for $q$, a natural approach to sampling from $\pi$ is to define a DDM approximation based on the pre-trained score approximations for $q_{k}$. Following the DDM approach, the basic idea here is to sample sequentially from the marginals $\pi_{k}(\boldsymbol{x}_{k})\vcentcolon=\int\pi(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\,\mathrm{d}\boldsymbol{x}_{0}$ by relying on approximations of the conditional scores $(\nabla\log\pi_{k})_{k=1}^{n}$. The latter can be expressed in terms of the unconditional scores by using
$\pi_{k}(\boldsymbol{x}_{k})\propto\int p({\boldsymbol{y}}|\boldsymbol{x}_{0})q(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\,\mathrm{d}\boldsymbol{x}_{0}\,,$ (2.2)
where after defining the backward transition kernel $q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})\vcentcolon=q(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})\big{/}q_{k}(\boldsymbol{x}_{k})$ yields
$\pi_{k}(\boldsymbol{x}_{k})\propto p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})q_{k}(\boldsymbol{x}_{k})\,,\quad\text{where}\quad p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})\vcentcolon=\int p({\boldsymbol{y}}|\boldsymbol{x}_{0})q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})\,\mathrm{d}\boldsymbol{x}_{0}\,.$ (2.3)
@@
$\nabla\log\pi_{k}(\boldsymbol{x}_{k})=\nabla\log q_{k}(\boldsymbol{x}_{k})+\nabla\log p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})\,.$
While we have a parametric approximation of the first term on the r.h.s., the bottleneck is the second term, which is intractable due to the integration of $p({\boldsymbol{y}}|\cdot)$ against the conditional distribution $\smash{q_{0|k}(\cdot|\boldsymbol{x}_{k})}$.
@@
To circumvent this computational bottleneck, previous works have proposed approximations that involve replacing $q_{0|k}(\cdot|\boldsymbol{x}_{k})$ in (2.3) with either a Dirac delta mass Chung et al. (2023) or a Gaussian approximation Ho et al. (2022); Song et al. (2023a)
$\smash{q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})\approx\mathcal{N}(\boldsymbol{x}_{0};\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k}),v_{0|k}\mathbf{I}_{d})}\,,$ (2.4)
@@
where $\smash{\boldsymbol{m}^{\theta}_{0|k}}$ is a parametric approximation of $\smash{\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k})\vcentcolon=\int\boldsymbol{x}_{0}\,q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})\,\mathrm{d}\boldsymbol{x}_{0}}$ and $v_{0|k}$ is a tuning parameter. Using Tweedie’s formula Robbins (1956), it holds that $\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k})=(-\boldsymbol{x}_{k}+\sqrt{\alpha_{k}}\,\nabla\log q_{k}(\boldsymbol{x}_{k}))/v_{k}$, which suggests the approximation $\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k})\vcentcolon=(-\boldsymbol{x}_{k}+\sqrt{\alpha_{k}}\boldsymbol{s}^{\theta}_{k}(\boldsymbol{x}_{k}))/v_{k}\,.$ As for the variance parameter $v_{0|k}$, Ho et al. (2022) suggest setting $v_{0|k}=v_{k}/\alpha_{k}$, while Song et al. (2023a) use $v_{0|k}=v_{k}$. When $p({\boldsymbol{y}}|\cdot)$ is the likelihood of a linear model, i.e., $p({\boldsymbol{y}}|\boldsymbol{x})\vcentcolon=\mathcal{N}({\boldsymbol{y}};\mathbf{A}\boldsymbol{x},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}})$, for $\mathbf{A}\in\mathbb{R}^{{d_{\boldsymbol{y}}}\times{d}}$ and $\sigma_{\boldsymbol{y}}>0$, it can be exactly integrated against the Gaussian approximation (2.4), yielding $p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})\approx\mathcal{N}({\boldsymbol{y}};\mathbf{A}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k}),\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}}+v_{0|k}\mathbf{A}\mathbf{A}^{\intercal})$. Chung et al. (2023), on the other hand, use the pointwise approximation, yielding $p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})\approx p({\boldsymbol{y}}|\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k}))$. We denote by $\tilde{p}_{k}({\boldsymbol{y}}|\cdot)$ the resulting approximation stemming from any of these methods. Approximate samples from the posterior distribution are then drawn by simulating a backward Markov chain $(\mathbf{X}_{k})_{k=n}^{0}$, where $\mathbf{X}_{n}\sim p_{n}$ and then, given $\mathbf{X}_{k+1}$, $\mathbf{X}_{k}$ is obtained via the update
@@
$\mathbf{\mathbf{X}}_{k}\vcentcolon=\tilde{\mathbf{X}}_{k}+w_{k+1}(\mathbf{X}_{k+1})\nabla\log\tilde{p}_{k+1}({\boldsymbol{y}}|\mathbf{X}_{k+1})\,,\quad\text{where}\quad\tilde{\mathbf{X}}_{k}\sim p^{\theta}_{k|k+1}(\cdot|\mathbf{X}_{k+1})\,,$ (2.5)
@@
and $w_{k+1}$ is a method-dependent weighting function. For instance, Chung et al. (2023) use $w_{k+1}(\mathbf{X}_{k+1})=\zeta\sigma_{\boldsymbol{y}}^{2}/\|{\boldsymbol{y}}-\mathbf{A}\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\mathbf{X}_{k+1})\|$, with $\zeta\in[0,1]$. We note that the update (2.5) in general involves a vector-Jacobian product, e.g., considering the approximation suggested by Chung et al. (2023), $\nabla_{\boldsymbol{x}_{k}}\log\tilde{p}_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})=\nabla_{\boldsymbol{x}_{k}}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k})^{\intercal}\nabla_{\boldsymbol{x}_{0}}\left.\log p({\boldsymbol{y}}|\boldsymbol{x}_{0})\right|_{\boldsymbol{x}_{0}=\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k})}$, which incurs an additional computational cost compared to an unconditional diffusion step.